40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 26 19:35:16 2019, Jon, Update, SUS, New c1susaux installed, passed first round of scripted testing IMG_3254.jpgIMG_3256.jpg
    Reply  Fri May 3 12:48:40 2019, gautam, Update, SUS, 1X4/1X5 cable admin D38CC485-1EB6-4B34-9EB1-2CB1E809A21A.jpeg
       Reply  Tue Jul 2 12:30:53 2019, gautam, Update, Electronics, Acromag crate switched to Sorensens 1X5Sorensens.pdfCDS_20190702.png
          Reply  Tue Jul 2 16:57:09 2019, gautam, Update, CDS, c1sus is flaky c1sus_timing.png
Message ID: 14592     Entry time: Fri May 3 12:48:40 2019     In reply to: 14581     Reply to this: 14718
Author: gautam 
Type: Update 
Category: SUS 
Subject: 1X4/1X5 cable admin 

Chub and I crossed off some of these items today morning. The last bullet was addressed by Jon yesterday. I added a couple of new bullets.

The new power connectors will arrive next week, at which point we will install them. Note that there is no 24V Sorensen available, only 20V.

I am running a test on the 2W Mephisto for which I wanted the diagnostics connector plugged in again and Acromag channels to record them. So we set up the highly non-ideal but temporary set up shown in Attachment #1. This will be cleaned up by Monday evening latest.

update 1630 Monday 5/6: the sketchy PSL acromag setup has been disassembled.

Quote:
 
  • Take photos of the new setup, cabling.
  • Remove the old c1susaux crate from the rack to free up space, possibly put the PSL monitoring acromag chassis there.
  • Test that the OSEM PD whitening switching is working for all 8 vertex optics.(verified as of 5/3/19 5pm)
  • New 15V and 24V power cables with standard LIGO connectors need to be run from the Sorensenn supplies in 1X5. The chassis is currently powered by bench supplies sitting on a cart behind the rack.
  • All 24 new DB-37 signal cables need to be labeled.
  • New 96-pin DIN connectors need to be put on two ribbon cables (1Y5_80 B, 1Y5_81) in the 1X4 rack. We had to break these connectors to remove them from the back of the eurcrates.
  • General cleanup of any cables, etc. left around the rack. We cleaned up most things this evening.
  • Rename the host computer c1susaux2 --> c1susaux, and update the DNS lookup tables on chiara.
Attachment 1: D38CC485-1EB6-4B34-9EB1-2CB1E809A21A.jpeg  1.737 MB  Uploaded Fri May 3 13:59:52 2019  | Show | Show all
ELOG V3.1.3-